Suppose P1(Q) and P2(Q) are demand and supply curves, respectively, where P1(Q) = 200e^−0.01Q and P2(Q) = 20e^0.08Q and where price P is in dollars for a quantity Q. Suppose there is a quantity regulation (quota) at 10 units.

(A) How much consumer surplus is lost from the effects of this regulation?
Answer #1

Loss of consumer surplus= Area A+ Area B.

Area A= (180.967-154.853)*10=$ 261.14

Area B= 0.5*(180.967-154.853)(25.584-10)= 0.5*26.114*15.584= $203.48

Loss of consumer surplus=$ 464.62